OpenStudy (anonymous):

Please help me! Medal+Fanning! Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). (5 points) Part B: How much time does Ted spend on practicing the breast stroke every day? (3 points)

OpenStudy (doc.brown):

What's x?

OpenStudy (anonymous):

Butterfly stroke every day is x and breast stroke everyday is y

OpenStudy (doc.brown):

Does x = y?

OpenStudy (anonymous):

I don't think so...

OpenStudy (doc.brown):

What's x + y?

OpenStudy (anonymous):

It would equal 50

OpenStudy (doc.brown):

Please write it as a linear equation.

OpenStudy (anonymous):

x+y=50

OpenStudy (doc.brown):

What does he do longer, breast stroke or butterfly stroke?

OpenStudy (anonymous):

Breast stroke.

OpenStudy (doc.brown):

How much longer?

OpenStudy (anonymous):

20 minutes longer

OpenStudy (doc.brown):

You said breast stroke was y and it's 20 bigger than x... so what's x?

OpenStudy (anonymous):

x is butterfly stoke

OpenStudy (doc.brown):

You said \(x\neq y\) because y is 20 larger.

OpenStudy (anonymous):

Yeah. :)

OpenStudy (doc.brown):

so\[x=y-20\]

OpenStudy (doc.brown):

What's y, you ask?\[x=y-20\]\[x+20=y-20+20\]\[x+20=y+0\]\[y=x+20\]

OpenStudy (anonymous):

Oh, okay, that makes soo much sense, thank u!

OpenStudy (doc.brown):

Here you take the time he does the butterfly stroke (x) and add twenty minutes.

OpenStudy (doc.brown):

Part A, two linear equations\[x=y-20\]\[y=x+20\]

OpenStudy (doc.brown):

Part B You know two things\[x+y=50\]and\[x=y-20\]Solve for y.

OpenStudy (anonymous):

Okay...how would you solve that? I'm confused.

OpenStudy (doc.brown):

Stick this\[x=\color{red}{y-20}\]in this\[\color{red}{x}+y=50\]

OpenStudy (anonymous):

So it would be like.... x(y-20)+y=50

OpenStudy (doc.brown):

\[x=y-20\]x equals y minus 20 so you need to replace the x with something that is equal to x\[\color{red}{y-20}+y=50\]

OpenStudy (anonymous):

Oh, okay, thanks!

OpenStudy (anonymous):

Then, what?

OpenStudy (doc.brown):

Add twenty to both sides.

OpenStudy (anonymous):

y-20+y=50 +20 +20 y-y=70

OpenStudy (doc.brown):

Watch the sign, you have two positive y's

OpenStudy (anonymous):

Oh... so... 2y=70

OpenStudy (doc.brown):

divide both sides by 2

OpenStudy (anonymous):

so y=35?

OpenStudy (doc.brown):

Check it. If y=35 and x=y-20

OpenStudy (anonymous):

35-20=15

OpenStudy (doc.brown):

So if he spends 15 minutes doing butterfly and 35 minutes doing breast... Does that sound reasonable?

OpenStudy (anonymous):

Yeah, that equals 50! :) Thanks soo much!

OpenStudy (doc.brown):

Well done.

OpenStudy (anonymous):

Thanks for ur help!